### v2

# MedPix 2.0: A Comprehensive Multimodal Biomedical Data set for Advanced AI Applications with Retrieval Augmented Generation and Knowledge Graphs

Abstract
The increasing interest in developing Artificial Intelligence applications in the medical domain, suffers from the lack of high-quality data set, mainly due to privacy-related issues. In addition, the recent increase in large multimodal models (LMM) leads to the need for multimodal medical data sets, where clinical reports and findings are attached to the corresponding CT or MRI scans. This paper illustrates the entire workflow for building the MedPix 2.0 data set. Starting with the well-known multimodal data set MedPix®, mainly used by physicians, nurses, and healthcare students for Continuing Medical Education purposes, a semi-automatic pipeline was developed to extract visual and textual data followed by a manual curing procedure in which noisy samples were removed, thus creating a MongoDB database. Along with the data set, we developed a GUI aimed at navigating efficiently the MongoDB instance and obtaining the raw data that can be easily used for training and/or fine-tuning LMMs. To enforce this point, in this work, we first recall DR-Minerva, a RAG-based LMM trained using MedPix 2.0. DR-Minerva predicts the body part and the modality used to scan its input image. We also propose the extension of DR-Minerva with a Knowledge Graph that uses Llama 3.1 Instruct 8B, and leverages MedPix 2.0. The resulting architecture can be queried in a end-to-end manner, as a medical decision support system. MedPix 2.0 is available on GitHub.
Keywords:
MedPix MongoDB Biomedical Dataset MLLM CLIP Decision Support System
1 Introduction
The rise of computer-based applications in recent years has strongly favored the digitization of historically analogous processes in the management and analysis of biomedical data. In turn, the emergence of technologies based on Artificial Intelligence (AI) prompted the development of increasingly precise models to support diagnosis for the construction of personalized treatments. In the biomedical domain, in fact, different subareas can benefit from AI-based systems, ranging from the administrative field (queue management in emergency areas, centralized integration of EHRs, and so on) to the clinical one, thanks to the use of AI that can efficiently extract useful features to reliably achieve diagnosis.
One of the fundamental requirements of these applications is their trustworthiness, as they must help physicians with a high degree of confidence, being able to provide reliable predictions and classifications. However, AI models require a considerable amount of data to achieve good performance. Data availability makes expansion into the biomedical domain much more complex than in other ones. One of the main problems lies in the availability of data sets to allow the scientific community to develop new AI approaches. This problem arises mainly from the sensitive nature of data that encompass privacy issues, and this makes it difficult to build public data sets containing images and/or clinical reports to be available to the scientific community.
To overcome these obstacles, the European Community created the European Health Data Space (EHDS) [^1] https://www.european-health-data-space.com/. This is a health-specific ecosystem composed of common rules, standards and practices, infrastructure, and a governance framework that aims to empower people through increased digital access and control of their personal electronic health data. The EHDS promotes a single market for electronic health record systems, relevant medical devices, and high-risk artificial intelligence systems. Finally, the EHDS aims at providing the researchers, which make use of health data with a trusted framework to control the entire analytical process [26, 31]. The systemic approach proposed by the EHDS will provide a controlled pool of both shared data and applications. It will allow AI in the medical field to access certified and controlled data by overcoming both medical and engineering obstacles through new solid foundations on which a new generation of AI-based health applications can rely on.
In view of the implications provided by the EHDS on the application side, researchers wishing to pursue their activity in the biomedical domain by developing Large Multimodal Models (LMMs) should be able to find and optimize the data sets that are available, in a way that maximizes the available public resources. To date, there are not so many data sets containing both images (CT and/or MRI) and medical reports. One of the most well-known data sets is MedPix® [^2] https://medpix.nlm.nih.gov/home, a free open-access online database of medical images, teaching cases, and clinical topics, integrating images and textual metadata. MedPix® includes more than 12,000 patient case scenarios, 9,000 topics, and nearly 59,000 images. The contribution proposed in this paper lies in both curation and a new arrangement of MedPix® data in a non-relational database based on MongoDB. This new arrangement of the MedPix® structure can be regarded as a brand new data set that we called MedPix 2.0. MedPix 2.0 makes original data accessible and well structured, as it is very easy to build views capable of creating ready-made subsets for training LMMs. Database queries have been further simplified by developing a user-friendly GUI. To enforce usability of MedPix 2.0, the user is made able to pose her/his query, and browsing the results in very close way as in the original website. Another contribution of this work is a direct demonstration of how MedPix 2.0 makes it easy to train and test LMMs. At first, we recall the architecture and the training phase of DR-Minerva [29], a RAG-based LMM developed by the authors that leverages the Minerva LLM [23] and uses Flamingo [1] for multi-modality. We used our MongoDB data source to generate the train/validation/test split used for DR-Minerva.
Then, we propose a Knowledge Graph (KG) built from MedPix 2.0, using Llama 3.1 Instruct 8B [20]. The KG can be queried to obtain information about medical diseases and their diagnoses. The entire system, that is the KG coupled with DR-Minerva, can be queried end-to-end starting from a medical image to obtain suggestions about the most probable diagnosis with a free-text answer, thus working as a medical decision support system.
The paper is arranged as follows: Section 2 illustrates the State Of The Art (SOTA) for medical multimodal data sets. The details on the building process of the curated data set are reported in Section 3 along with the implementation of the MongoDB database and the GUI. The experimental results and The architecture of DR-Minerva and the creation of the KG are reported and discussed in Section 4, while Section 5 illustrates the experimental results. Concluding remarks and future works are drawn in Section 6.
2 Related Works
Medical data sets for AI applications suffer from diverse problems, related both with the data and the peculiarity of the domain. First, there is a privacy issue, since clinical data contain private information about the patient, thus the process of creation of such type of data set has to start with an anonymization phase. To overcome this problem, researchers rely on either open-access and textbook data or they collaborate with hospitals to create data sets. The first two methods allow for large scalability, relying on anonymous data. On the other hand, anonymization must be carried out from scratch when dealing with hospitals. Moreover, multimodal medical data suffer from scarcity when compared with other data sets related to different domains such as MS-COCO [16].
One of the most used open-access multi modal database for developing medical data sets is PubMed Central® (PMC) [^3] https://www.ncbi.nlm.nih.gov/pmc/. This is a widely used free archive of biomedical scientific literature, from which it is possible to build one’s own data sets via semi-automatic procedures. In PMC, data are anonymized, and high-quality captions can be extracted from the medical research papers the images belong to. The following multimodal data sets were extracted from PMC: ROCO [25], MedICaT [30], PMC-OA [17]. ROCO contains pairs of radiology images and the corresponding captions, and it incorporates an out-of-class set to improve prediction and classification performances. MedICaT is a disjoint data set from ROCO that is mainly composed by radiology images and provides manually annotations for sub-figures. PMC-OA is the larger than the previous ones, and it keeps a variety of diagnostic procedures, diseases, and findings, while introducing sub-figure separation.
VQA-RAD [13] is a data set derived from MedPix®, and it collects a subset of radiological images, while providing Question-Answer (QA) pairs validated by domain experts.
Another source of available high-quality data are textbooks: PathVQA [8] is a Visual Question Answering (VQA) data set that collects both closed- and open-ended QA pairs, which are extracted from both pathology textbooks and online digital libraries via a semi-automated pipeline.
On the other side, open access data sets like MIMIC-CXR [11], IU-Xray [5] and SLAKE [18] are manually annotated by domain experts. Both MINIC-CXR and IU-Xray are chest radiography data sets derived from hospital’s clinical cases. Both if them contain a semi-structured radiology report, describing the radiological findings of the images it is related to. SLAKE, on the contrary, collects images from different radiology open-access data sets and provides manual annotations and QA pairs given by experienced doctors in English and Chinese.
INSPECT [9] is a worth mentioning multimodal dataset that collects computed tomography pulmonary angiography (CTPA), radiology reports and structured data from EHRS, for pulmonary embolism diagnosis.
The aforementioned data sets are multimodal ones, and they are deeply focused on VQA tasks. Also some specialized data sets exist that provide only visual data, as UniToChest [4] or free-form textual data [28]. Other works worth mentioning are UniToBrain [7] that integrates a technical report with the scanning modality, and the E3C corpus [21] that is a multilingual data set of clinical reports.
Despite the proposed MedPix 2.0 data set may be close to VQA-RAD, they differ from both the sampling strategy and the samples themselves. In Med Pix 2.0, we do not integrate QA pairs, but different QA pairs can be created relying on the structured textual information provided in the data set itself, following the structure of the other one. Thus, the creation of more complex tasks is possible, such as document summarization or understanding.
With respect to the existing multimodal data sets, MedPix 2.0:
  1. 1.
    is derived from a freely open-access source, and it has no privacy-related issues;
  2. 2.
    offers a balanced variety of CT and MRI scans of different body parts;
  3. 3.
    for each image, a complete structured clinical case is provided.
Thanks to the annotation scheme we selected for the JSON documents in MedPix 2.0, the last point makes it suited for various tasks not only limited to document-level retrieval. Unfortunately, images are provided in PNG format thus limiting visual processing with respect to raw images in DICOM format.
3 MedPix 2.0
MedPix® is a free open-access multimodal online database of medical images, teaching cases, and clinical topics, managed by the National Library of Medicine (NLM) of the National Institutes of Health (NIH). It mainly serves as a support system for Continuing Medical Education (CME) of physicians, nurses, and healthcare students. The database collects clinical cases related to more than 12,000 patients. Each case contains at least one medical image, and the corresponding findings, discussion notes, diagnosis, differential diagnosis, treatment, and follow up. Textual information is reported in a semi-structured format. Attached to the clinical case, there is the topic section, where the disease under investigation is discussed in detail from an academic and general perspective.
In Fig. 1 an example from the MedPix® website is reported.
Figure 1: An example from the MedPix® web site where the original screenshots have been rearranged for visualizing purposes, and the colored labels have been added for clarity.
Despite the richness of the data set, its free availability, the possibility to add new cases, and the access to clinical cases of interest using wither keywords, the body part or the disease, it is not possible to access the raw data. This feature limits the usage of MedPix® for training multimodal AI systems. Therefore we decided to create a brand new structured version of this data set because it represents a high quality source for AI-based medical applications. MedPix 2.0 has been built essentially as a MongoDB instance that is released along with a suitable GUI aimed both at general purpose querying and extracting training data for AI models. Referring to the Fig.1, a MongoDB version of the data set was built using a semi-automated pipeline to create two kinds of JSON documents: the one collecting the information falling into the screenshots labeled as DESCRIPTION, and the one that gathers the information falling into the screenshots labeled both as CASE and TOPIC.
3.1 Data set extraction
Figure 2: Among the manually removed images there are presentation’s slide (a), manually (b) and automatic annotated images from DICOM viewer (c)
We decided to focus on a part of MedPix® that involves cases related with two diagnostic modalities, namely Computed Tomography (CT), and Magnetic Resonance Imaging (MRI). First of all, the images in the considered split were downloaded via Open-i®  [^4] https://openi.nlm.nih.gov/s: among them, there where some noisy images such as teaching materials or annotated images (Fig. 2) that are inconsistent as input image for training a LMM for which we expect a clear CT or MRI. A subsequent automatic scraping pipeline was implemented to extract the textual data related to the selected images using Selenium [^5] https://www.selenium.dev/ and Beautiful Soup [^6] https://www.crummy.com/software/BeautifulSoup/bs4/doc/. Finally, two kinds of JSON documents were devised to store, respectively, the information strictly connected with the images (descriptions document) and the one related to a clinical case (case-topic document). A one-to-many relation has been created between a clinical cases and images by embedding the U_id defined for a case-topic document in each descriptions document attached to each image related to the clinical case itself. An example of the two kinds of JSON documents is shown in Fig. 3.
Figure 3: An example of the JSON documents created in MedPix 2.0: (a) descriptions document, (b) case-topic document.
3.2 MongoDB representation
In order to process properly the data in MedPix 2.0 we built a MongoDB database to host all the JSON documents along with the images. The architectural choice stems not only from the nature of the data in MedPix 2.0 but also from considerations related to its high flexibility and scalability to distributed environments where also private multimodal medical data could be added to the original collections with the constraint of not being moved away from their generation site as it is the case of hospital generated information.
We built a MongoDB instance made by two collections, namely Image_ Descriptions, containing the descriptions documents, and Clinical_reports which contains all the case-topic documents. In our implementation, the images are stored in a separated folder, and they are accessed using a proper file:// URL built starting from their U_id. Finally, a view called Image_ Reports allow a direct access to both collections via their U_id.
A user-friendly GUI was built using PyQt5777https://www.riverbankcomputing.com/software/pyqt/ to allow an easy access to the database, querying it to obtain the desired data for visualization and/or download. As shown in Fig. 4.a, it is possible to select either the collection or the view to be queried and, add the query input in the relative fields. In Fig. 4.b, an example of the answer to a query is reported: samples matching the query are reported as a list of JSON objects and the user can save it or view a specific clinical case and/or image selected in list of the query results. An example of this view is shown in Fig. 5, where a MedPix® like GUI is reproduced to enforce usability for the users of the original website. In our GUI, the curated images scraped along with the texts are showed by default, but the user can choose to download the original (non curated) data related to the clinical case under investigation.
Figure 4: An example of the interface for querying the database where the user asks for CT scans of female patients (a) and the provided output (b).
Figure 5: An example of the view of a selected clinical case retrieved by the query posed in Fig. 4. The image, and the relative description are reported along with the information of the clinical case the image belongs to.
MedPix 2.0 and its querying interface can be a valid tool for both physicians and AI researchers since the desired data can be easily downloaded for further application, like training Deep Learning (DL) models where a large amount of structured data are required. The structured output downloaded from the interface, can be used in turn to fill a fixed template, and to provide a LMM with a rich textual prompt. Also multimodal tasks can be addressed if the textual descriptions are combined with the corresponding image, as we will demonstrate in Section 4. Our re-arranged version of MedPix®, contains the same information of the original dataset, but allows a easier usage for DL application in text-only, image-only and multimodal setup. This flexible structure is also suitable for adding new samples that meet the characteristics of the dataset, i.e. they should be a clear CT or MR scan with the associated textual data as in Fig 3.
4 Training a LMM using MedPix 2.0
This section reports an application scenario for using Medpix 2.0. Specifically, a training pipeline for a multimodal model was implemented and, starting from the images, through a RAG-based approach, textual and visual embedding are generated to predict the scan modality and the body part shown in the input image. The output at this stage is furthermore used to query a generative Large Language Model (LLM) which will generate a consistent answer for the end user through the use of a Knowledge Graph (KG), The answer contains a diagnostic suggestion, leveraging the acquired knowledge on both scan modality and body part and the knowledge in the KG.
4.1 Data preparation
The training phase of the model was conducted using MedPix 2.0, carefully divided into three splits: Train, Validation, and Test, with an 80:10:10 ratio. Correct stratification of the data required considerable effort. In fact, all the images referring to the same clinical case are reported in the same split, in order to reduce data leakage. The data obtained from this division are shown in Table 1 and 2.
Dataset #CT #MR Clinical Case
Train 878 775 535
Validation 84 113 67
Test 100 100 69
Total 1062 988 671
Table 1: Summary of images and clinical cases for each split.
Dataset Abdomen Head RUS* Thorax SaM**
Train 264 742 127 263 257
Validation 23 66 20 30 58
Test 32 76 11 41 40
Total 319 884 158 334 355
Table 2: Summary of images divided per macro area of the body for each split. (*RUS: Reproductive and Urinary System; ** SaM: Spine and Muscles)
4.2 RAG-based Flamingo
Figure 6 illustrates the overall RAG architecture used to test MedPix 2.0, which is called DR-Minerva [29]. DR-Minerva relies on Flamingo [1] and Minerva [23], a new LLM trained from scratch on English and Italian data as part of the activities in the PNRR FAIR Transversal Project 2: “Vision, Language, and Multimodal Challenges” [^8] https://fondazione-fair.it/en/transversal-projects/tp2-vision-language-and-multimodal-challenges/. FAIR TP2 is a collaborative research carried out by more than twenty Italian universities. The authors are engaged in developing LMMs for the biomedical domain. We chose Minerva because it is totally open. Not only the weights, but also the architecture is free along with the training data. Furthermore, its training set is half English and half Italian, making it suitable for future experimentation using Italian data directly.
Figure 6: Overview of DR-Minerva architecture.
In our implementation, we used Open Flamingo [2] in the 3B parameter version [^9] https://huggingface.co/openflamingo/OpenFlamingo-3B-vitl-mpt1b. The Flamingo architecture is well known for its strong in-context learning capabilities, making it adaptable to various domains [1]. We used the CLIP ViT-L/14 [27] as visual encoder and Minerva-3B  [^10] https://huggingface.co/sapienzanlp/Minerva-3B-base-v1.0 as language encoder. The table 3 shows the prompt used to query the model in order to obtain the desired response, which consists of both the scanning modality and the body part shown in the input image.
We developed a suitable RAG component for DR-Minerva [14] to improve the overall performance of the model at inference time, due to a prompt enriched with relevant information for the query. Because AI models must be as precise as possible, especially in the medical sector, we queried DR-Minerva with both the target medical image and a template made up of some personal information about the patient (e.g. age and gender), followed by the patient’s history. The most relevant clinical cases are then retrieved and added to the prompt in the form of few-shot learning examples, based on the patient’s history. The RAG component selects (up to) the four closest clinical cases to the query. The vector database in the RAG component has been filled suitably with textual information from clinical cases, which belong to the train split, and involve the age and gender of the patient, the relative clinical history, and the doctor’s diagnosis.
Our RAG is built using the LangChain framework [^11] https://www.langchain.com/, which in turn employs a FAISS vector database [6], where the data are saved using Linq-Embed-Mistral [12]. This model is regarded as the best in the Massive Text Embedding Benchmark (MTEB) [22] for Information Retrieval121212as in https://huggingface.co/spaces/mteb/leaderboard in June 2024.
Both text-encoding and text generation are demanded to Minerva LLM, a transformer decoder-only LLM ([32]) built upon the Mistral architecture ([10]) which was trained from scratch on 660B tokens, equally balanced between English and Italian. Flamingo leverages a perceiver-resampler network to extract visual tokens from the visual encoding. In turn, visual tokens are employed to condition the textual output generation by inserting cross-attention layers between the existing pretrained ones and frozen LLM layers [1].
Prompt Response
Given the following medical images and the patient
history, provide information about the scanning
modality and the body part shown in the image.
[…]
[… context … ]
[…]
[… context …]
[…]
The image is a
[…]
scan showing a
[…]
Table 3: The structure of the prompt is reported as well as the template of the corresponding expected answers
4.3 Knowledge Graph definition
MedPix 2.0 is actually a huge archive of historical clinical cases that can be exploited as external knowledge for sophisticated textual predictions. In addition to the information related to the particular case reported in the Case section, such as Findings and Discussion, there is also more general information related to the disease, like Disease Discussion in the Topic section (Fig. 3). Due to the inherent domain-specific semantics of this kind of textual information, classical RAG relying on vector databases is not efficient for precise retrieval. As a consequence, we decided to build a Knowledge Graph, which relies on both case-specific and general medical information that can be retrieved at inference time to generate a diagnostic suggestion for the provided input. The whole process for creating the KG and inferring suggestions is depicted in Fig. 7
Figure 7: Overview of the overall process for KG creation and inference.
To allow efficient relation extraction for building the KG, clinical cases belonging to the training set were arranged as documents according to the following template:
[…]
We used LlamaIndex [19] to generate the Knowledge Graph. In particular, the KG is built starting from document chunks whose fixed size is 4096 bytes. At most n triplets were extracted from each chunk to form the nodes and edges of the KG. Triplets were extracted using the default LlamaIndex query posed to Llama 3.1 8B Instruct [20]. This choice was due to the native multilingual support of the model and its instruction-following capabilities. We created six different KGs according to the maximum number of triplets to be extracted from each chunk (3, 5, 10) and the option to process the chunk in lowercase or not. The hyperparameter setups are reported in Table 4.
Setup name max rel. per chunk Chunks in lower-case
KG-s1 10 no
KG-s2 10 yes
KG-s3 3 no
KG-s4 3 yes
KG-s5 5 no
KG-s6 5 yes
Table 4: Overview of the hyperparameter setups used to generate the different versions of the KG.
Figure 8: Overview of the overall generated KG-s1 (a) and of its subgraph generated from the node ’Pulmonary involvement’.
The entire process of creation and inference over the KGs was run on a cluster with 2 NVIDIA A100 64 GB GPUs.
5 Results
The main contribution of this work, as regards the effectiveness of using MedPix 2.0 to train and test a complex multimodal pipeline, is the end-to-end LMM used for medical diagnosis support. The architecture of this system is depicted in Fig. 9, and couples the original DR-Minerva model with the KG built as reported in Section 4.3.
Figure 9: Overview of the generated architecture with RAG-based Flamingo and KG creation and inference.
Performance evaluation was carried out by querying all the KGs with different hyperparameter choices, using prompts of increasing complexity. Specifically, three prompts were constructed, which were called simple, mid-complex and complex respectively. All the prompts are reported as follows.
[…]
The three prompts use the information provided by DR-Minerva, which consists in the scanning modality and the body part shown in the input image. Prompts were designed to provide more and more information to the Knowledge Graph, as their complexity increases. The maximum complexity will be reached in the last prompt, where we want to emulate the query that a doctor might pose to a diagnostic support system. In this case, information concerning the patient’s clinical $\bigl{\langle}history\bigr{\rangle}$ is used in addition to the information extrapolated from the images supplied to the multimodal system.
The availability of a test set extrapolated from MedPix 2.0 allowed us to generate a golden answer according the following template.
[…]
Golden answers were used for the subsequent performance evaluation phase. Several metrics exist in the literature for automatic evaluation of LLM-generated text with respect to a set of reference or “golden” answers. These metrics score the degree of overlap between the reference answer $r=\{r_{1},\ldots,r_{m}\}$ and the hypothesis $h0\{h_{1},\ldots,h_{n}\}$ generated by the model. There are two approaches in this respect: character, word, or n-gram overlap and embedding similarity. BLEU [24] and ROUGE [15] belong to the first category. They are generally more tailored for evaluating Question-Answering and Machine Translation systems, for which a well-defined answer is known. Such measures require the generated text $h$ to adhere strictly to the golden answer $r$.
The generation of correct clinical text to support a doctor in the diagnostic process is a very different task. In this case, the same meaning can be expressed in a variety of forms that depend on many factors: the available knowledge about the particular case when the diagnosis is formulated, the different clinical practices in writing medical records, and so on. As a consequence, despite the generated golden answers and the dataset splits, we cannot expect the model to generate text exactly as the golden answer.
For these reasons, in our evaluation setup we preferred to use embedding similarity scores like METEOR [3] and BERT score [33]. In this case the generic $r_{j}$ and $h_{i}$ are embeddings and not words or n-grams. In particular, we selected the BERT score that is a neural framework to compare the embeddings of the tokens in $r$ and $h$ in terms of classical Precision, Recall, and F1 measure. Equation 1 reports the BERTs score formulation.
$\begin{array}[]{lll}&\text{BERT-P}&=\frac{1}{n}\sum_{i=1}^{n}\underset{r_{j}}{\max}\ \text{Sim}(h_{i},r_{j})\\ \\ &\text{BERT-R}&=\frac{1}{m}\sum_{j=1}^{m}\underset{h_{i}}{\max}\ \text{Sim}(r_{j},h_{i})\\ \\ &\text{BERT-F1}&=2\cdot\frac{\text{BERT-P}\cdot\text{BERT-R}}{\text{BERT-P}+\text{BERT-R}}.\end{array}$ (1)
$Sim(r_{j},h_{i})$ is the cosine similarity between the $j$-th reference token embedding and the $i$-th hypothesis token embedding, respectively. Table 5 shows the results obtained using all the three prompts.
// Simple Mid Complex
KG-setup Location BERT-Precision BERT-Recall BERT-F1 BERT-Precision BERT-Recall BERT-F1 BERT-Precision BERT-Recall BERT-F1
KG-s1 Head 0.8027 0.7840 0.7932 0.7870 0.7848 0.7857 0.7971 0.7858 0.7913
Thorax 0.8003 0.7801 0.79 0.7853 0.7806 0.7827 0.7875 0.7822 0.7847
SaM** 0.7888 0.7932 0.791 0.7964 0.7851 0.7905 0.7846 0.7885 0.7864
Abdomen 0.8008 0.7855 0.793 0.7859 0.7782 0.7818 0.7843 0.7817 0.7829
RUS* 0.8050 0.7886 0.7967 0.7774 0.7787 0.778 0.7884 0.7766 0.7823
KG-s2 Head 0.8090 0.7900 0.7994 0.7985 0.7863 0.7921 0.7888 0.7802 0.7843
Thorax 0.8039 0.7851 0.7943 0.7981 0.7746 0.7856 0.7873 0.7821 0.7846
SaM** 0.7981 0.7950 0.7964 0.7888 0.7908 0.7898 0.7949 0.7860 0.7903
Abdomen 0.7980 0.7845 0.7912 0.7882 0.7830 0.7853 0.7902 0.7863 0.7882
RUS* 0.7971 0.7850 0.791 0.8011 0.7817 0.7912 0.8063 0.7890 0.7975
KG-s3 Head 0.7990 0.7855 0.7921 0.7991 0.7825 0.7905 0.8019 0.7872 0.7944
Thorax 0.7987 0.7840 0.7912 0.7885 0.7800 0.784 0.7858 0.7856 0.7855
SaM** 0.8058 0.7958 0.8007 0.7820 0.7906 0.7862 0.7993 0.7902 0.7947
Abdomen 0.7960 0.7874 0.7916 0.7815 0.7742 0.7775 0.7844 0.7846 0.784
RUS* 0.8021 0.7799 0.7908 0.7972 0.7781 0.7874 0.7985 0.7878 0.7929
KG-s4 Head 0.7989 0.7799 0.7891 0.7915 0.7829 0.787 0.7952 0.7843 0.7895
Thorax 0.8031 0.7784 0.7905 0.7869 0.7764 0.7815 0.7922 0.7822 0.7868
SaM** 0.8023 0.7934 0.7977 0.7870 0.7874 0.7871 0.8006 0.7878 0.7939
Abdomen 0.7989 0.7769 0.7875 0.7975 0.7837 0.7905 0.7779 0.7765 0.7771
RUS* 0.7979 0.7819 0.7898 0.7753 0.7798 0.7774 0.7937 0.7819 0.7877
KG-s5 Head 0.7999 0.7858 0.7927 0.7960 0.7876 0.7917 0.7945 0.7814 0.7877
Thorax 0.7960 0.7808 0.7882 0.7915 0.7769 0.784 0.7883 0.7827 0.7852
SaM** 0.8037 0.7953 0.7995 0.7888 0.7932 0.7908 0.8049 0.7925 0.7985
Abdomen 0.8012 0.7772 0.7889 0.7840 0.7795 0.7816 0.8028 0.7820 0.7921
RUS* 0.7861 0.7860 0.786 0.7876 0.7871 0.7873 0.8025 0.7898 0.796
KG-s6 Head 0.7910 0.7864 0.7885 0.7923 0.7801 0.786 0.7931 0.7799 0.7862
Thorax 0.7859 0.7812 0.7835 0.7720 0.7754 0.7735 0.7879 0.7808 0.7842
SaM** 0.7832 0.7893 0.7861 0.7957 0.7893 0.7924 0.8106 0.7918 0.8009
Abdomen 0.7807 0.7816 0.7811 0.7958 0.7875 0.7916 0.7886 0.7801 0.7841
RUS* 0.7973 0.7826 0.7898 0.7849 0.7842 0.7845 0.7994 0.7886 0.7938
Table 5: Summary of BERT metrics calculated on all KG setups for each of the proposed prompts. Bold values indicate the best score for each KG and body part respectively
The results show that the prompt simple achieves best performance for the models using KG-s1 and KG-s2, that are the KGs with the highest number of relations per chunk. The prompt complex performs best for the model using KG-s6, and it is close to the prompt simple for the models using KG-s3. No particular prompt prevails over the others in the models using KG-s4 and KG-s5.
[…]
At inference time, Llama 3.1 8B Instruct is invoked for answer generation by retrieving relevant information from the KGs, leveraging the given question. We used a low temperature of $0.00001$, thus the generative capabilities of the model are limited, and then, it is forced to produce an answer as much close as possible to the retrieved information. Moreover, to prevent the model in generating loop, the no_repeat_ngram_size parameter was set to $2$. The query strategy over the KG was done using the tree_summarize response mode and a value equal to $5$ as for similarity_top_k parameter, that is the default parameters for the retriever. We report four end-to-end inferences using the better prompts and KGs, namely simple prompt and KG-s1 (Fig 10) and complex prompt and KG-s6 11). Both the image, the input prompt to the model, the generated answer and the BERT-F1 score are reported.
Figure 10: Two example of end-to-end queries to the entire system leveraging KG-s1 and simple prompt, along with inputs and the final generated answer.
Figure 11: Two example of end-to-end queries to the entire system leveraging KG-s6 and complex prompt, along with inputs and the final generated answer.
6 Conclusions and Future work
In this work we presented MedPix 2.0, a multimodal data set of clinical reports, CT and MR scans. We devised a semi-automated pipeline to download and curate the images in the original data sets, while structuring the textual information as a set of JSON document collections which had been used to build a proper MongoDB instance. The NoSQL version of the data set can be accessed and queried with a usable GUI that has been developed purposely. Using the GUI one can browse the data set in the same manner as in the original data set, and can download the structured output of the query that is suitable for training AI models. In fact, MedPix® and its MongoDB interface, represent in our view a relevant starting point for the development of AI multimodal models in the medical domain. such as Information Extraction systems tailored for clinical reports, automated analysis of the medical images, or Generative AI models for clinical report generation as part of a Medical Decision Support System. All these systems can rely on MedPix 2.0 as a structured source of data containing both clinical cases and medical explanations about the disease under investigation.
To demonstrate this point, we recalled the architecture of DR-Minerva, a RAG-based Flamingo architecture already developed by the authors, and presented a brand new LMM architecture that couples DR-Minerva with a Knowledge Graph created automatically relying on Llama 3.1 8B Instruct. Different KG structures were tested, starting from the MedPix 2.0 test set along with prompts of increasing complexity as regards the context provided to the model. The results we obtained are very satisfactory but we are actively working to improve the model generation ability by means of a structural and semi-automatic organization of the KG, relying on external medical ontologies.
The scalability of the developed MongoDB database, makes it suitable for future extensions with the possibility to add brand new clinical cases or existing cases from data set mentioned in Section 2, that have to be compliant with privacy regulations and follow the required information structure. Moreover, the inherent distributed nature of MongoDB allows for creating huge databases across different wards where the data owned by a single institution do not need to be explicitly moved out of the hospital thus violating privacy regulations. The structure of MedPix 2.0 could also serve as a guide to develop suitable connectors to share allowed data in the EHDS.
New cases can also be easily added to the KG, thus expanding the clinical knowledge base used by the system in the inference phase. Further improvement is being done on the GUI, providing the user with advanced data visualization tools like the possibility to interactively compare similar cases, thus helping physicians during the diagnostic phase.
{credits}
6.1 Data and Code Availability
6.1.1 Acknowledgements
This work is supported by the cup project B73C22000810001, project code ECS_00000022 “SAMOTHRACE” (Sicilian MicronanoTech Research And Innovation Center).
[…]
The authors declares that they have no relevant or material financial interests that relate to the research described in this paper.