### v4

# MedPix 2.0: A Comprehensive Multimodal Biomedical Data set for Advanced AI Applications

Abstract
The increasing interest in developing Artificial Intelligence applications in the medical domain, suffers from the lack of high-quality data set, mainly due to privacy-related issues. Moreover, the recent rising of Large Multimodal Models (LMM) leads to a need for multimodal medical data sets, where clinical reports and findings are attached to the corresponding CT or MR scans. This paper illustrates the entire workflow for building the data set MedPix 2.0. Starting from the well-known multimodal data set MedPix®, mainly used by physicians, nurses and healthcare students for Continuing Medical Education purposes, a semi-automatic pipeline was developed to extract visual and textual data followed by a manual curing procedure where noisy samples were removed, thus creating a MongoDB database. Along with the data set, we developed a GUI aimed at navigating efficiently the MongoDB instance, and obtaining the raw data that can be easily used for training and/or fine-tuning LMMs. To enforce this point, we also propose a CLIP-based model trained on MedPix 2.0 for scanning modality and location classification tasks. MedPix 2.0 is available on GitHub [^1] https://github.com/CHILab1/MedPix-2.0.
Keywords:
MedPix MongoDB Biomedical data set LMM CLIP Decision Support System
1 Introduction
The rise of computer-based applications in recent years has strongly favoured the digitisation of historically analogue processes in the management and analysis of biomedical data. In turn, the emergence of technologies based on Artificial Intelligence (AI) prompted the development of increasingly precise models to support diagnosis for the construction of personalised treatments. In the biomedical domain, in fact, different sub-areas can benefit of AI-based systems, ranging from the administrative field (e.g., better management of queues in emergency areas, centralised integration of medical records) to the clinical one, thanks to the use of AI that can efficiently extract useful features to reliably achieve diagnosis.
One of the fundamental requirements of these applications consists in their trustworthiness, which must help the physician with high confidence, being able to provide reliable predictions and classifications. AI models, however, require a considerable amount of data to achieve these results, making this expansion into the biomedical domain much more complex than it actually would be in a different domain. One of the main problems lies in the availability of data sets to allow the scientific community for developing new AI approaches. This problem arises mainly from the sensitive nature of the data that has to cope with privacy issues, and this makes it difficult to build public data sets containing images and/or clinical reports to be available for the scientific community.
In order to start to overcome these obstacles, the European Community has constituted the European Health Data Space (EHDS) [^2] https://www.european-health-data-space.com/. This is a health-specific ecosystem composed of common rules, standards and practices, infrastructure and a governance framework that aims to empower people through increased digital access and control of their personal electronic health data. The EHDS promotes a single market for electronic health record systems, relevant medical devices, and high-risk artificial intelligence systems. Finally, the EHDS aims at providing the research that makes use of health data with a trusted framework for controlling the whole analytical process [15, 19]. The systemic approach proposed by the EHDS will provide a controlled pool of both shared data and applications thus allowing AI in the medical field to access to certified and controlled data by overcoming both medical and engineering obstacles through new solid foundations on which a new generation of AI-based health applications can rely on.
In view of the implications provided by the EHDS on the application side, researchers wishing to pursue their research in the biomedical domain by developing Large Multimodal Models (LMMs) should be able to find and optimise the data sets that are available, in a way that maximises the available public resources. To date, the data sets containing both images (CT and/or MRI) and medical reports are not many. One of the most renowned data sets is MedPix® [^3] https://medpix.nlm.nih.gov/home, a free open-access online database of medical images, teaching cases, and clinical topics, integrating images and textual metadata. MedPix® includes more than 12,000 patient case scenarios, 9,000 topics, and nearly 59,000 images. The contribution proposed in this paper, lies in a new organisation of a non-relational database based on MongoDB, which reorganises the structure of MedPix®, making it accessible and structured, through the construction of views capable of creating ready-made subsets for training LMMs. Database queries are further simplified through the development of a user-friendly GUI which, allows the user to pose her/his query thus browsing the results in very close way as in the original website. In addition, we used our MongoDB data source to create a data set for training a CLIP-based model to classify an input medical image, providing information about both the scanning modality and the body part shown, thus showing a practical application and study case the data set can be used on.
The paper is arranged as follows: Section 2 illustrates State Of The Art (SOTA) for medical multimodal data sets, the details on the building process of the curated data set are reported in Section 3 along with the MongoDB implementation and the proposed GUI. The experimental results of a LMM trained over the presented data set, demonstrating a use-case of MedPix 2.0, are reported and discussed in Section 4. Concluding remarks and future works are drawn in Section 5.
2 Related Works
Medical data sets for AI applications suffer from diverse problems, related both with the data and the peculiarity of the domain. First, there is a privacy issue, since clinical data contain private information about the patient, thus the process of creation of such type of data set has to start with an anonymization phase. To overcome this problem, researchers rely on either open-access and textbook data or they collaborate with hospitals to create data sets. The first two methods allow for large scalability, relying on anonymous data. On the other hand, anonymization must be carried out from scratch when dealing with hospitals. Moreover, multimodal medical data suffer from scarcity when compared with other data sets related to different domains such as MS-COCO [10].
One of the most used open-access multi modal database for developing medical data sets is PubMed Central® (PMC) [^4] https://www.ncbi.nlm.nih.gov/pmc/. This is a widely used free archive of biomedical scientific literature, from which it is possible to build one’s own data sets via semi-automatic procedures. In PMC, data are anonymized, and high-quality captions can be extracted from the medical research papers the images belong to. The following multimodal data sets were extracted from PMC: ROCO [14], MedICaT [18], PMC-OA [11]. ROCO contains pairs of radiology images and the corresponding captions, and it incorporates an out-of-class set to improve prediction and classification performances. MedICaT is a disjoint data set from ROCO that is mainly composed by radiology images and provides manually annotations for sub-figures. PMC-OA is the larger than the previous ones, and it keeps a variety of diagnostic procedures, diseases, and findings, while introducing sub-figure separation.
VQA-RAD [8] is a data set derived from MedPix®, and it collects a subset of radiological images, while providing Question-Answer (QA) pairs validated by domain experts.
Another source of available high-quality data are textbooks: PathVQA [4] is a Visual Question Answering (VQA) data set that collects both closed- and open-ended QA pairs, which are extracted from both pathology textbooks and online digital libraries via a semi-automated pipeline.
On the other side, open access data sets like MIMIC-CXR [6], IU-Xray [2] and SLAKE [12] are manually annotated by domain experts. Both MINIC-CXR and IU-Xray are chest radiography data sets derived from hospital’s clinical cases. Both if them contain a semi-structured radiology report, describing the radiological findings of the images it is related to. SLAKE, on the contrary, collects images from different radiology open-access data sets and provides manual annotations and QA pairs given by experienced doctors in English and Chinese.
INSPECT [5] is a worth mentioning multimodal dataset that collects computed tomography pulmonary angiography (CTPA), radiology reports and structured data from EHRS, for pulmonary embolism diagnosis.
The aforementioned data sets are multimodal ones, and they are deeply focused on VQA tasks. Also some specialized data sets exist that provide only visual data, as UniToChest [1] or free-form textual data [17]. Other works worth mentioning are UniToBrain [3] that integrates a technical report with the scanning modality, and the E3C corpus [13] that is a multilingual data set of clinical reports.
Despite the proposed MedPix 2.0 data set may be close to VQA-RAD, they differ from both the sampling strategy and the samples themselves. In Med Pix 2.0, we do not integrate QA pairs, but different QA pairs can be created relying on the structured textual information provided in the data set itself, following the structure of the other one. Thus, the creation of more complex tasks is possible, such as document summarization or understanding.
With respect to the existing multimodal data sets, MedPix 2.0:
  1. 1.
    is derived from a freely open-access source, and it has no privacy-related issues;
  2. 2.
    offers a balanced variety of CT and MRI scans of different body parts;
  3. 3.
    for each image, a complete structured clinical case is provided.
Thanks to the annotation scheme we selected for the JSON documents in MedPix 2.0, the last point makes it suited for various tasks not only limited to document-level retrieval. Unfortunately, images are provided in PNG format thus limiting visual processing with respect to raw images in DICOM format.
3 MedPix 2.0
MedPix® is a free open-access multimodal online database of medical images, teaching cases, and clinical topics, managed by the National Library of Medicine (NLM) of the National Institutes of Health (NIH). It mainly serves as a support system for Continuing Medical Education (CME) of physicians, nurses, and healthcare students. The database collects clinical cases related to more than 12,000 patients. Each case contains at least one medical image, and the corresponding findings, discussion notes, diagnosis, differential diagnosis, treatment, and follow up. Textual information is reported in a semi-structured format. Attached to the clinical case, there is the topic section, where the disease under investigation is discussed in detail from an academic and general perspective.
In Fig. 1 an example from the MedPix® website is reported.
Figure 1: An example from the MedPix® web site where the original screenshots have been rearranged for visualizing purposes, and the colored labels have been added for clarity.
Despite the richness of the data set, its free availability, the possibility to add new cases, and the access to clinical cases of interest using wither keywords, the body part or the disease, it is not possible to access to the raw data. This feature limits the usage of MedPix® for training multimodal AI systems. Therefore we decided to create a brand new structured version of this data set because it represents a high quality source for AI-based medical applications. MedPix 2.0 has been built essentially as a MongoDB instance that is released along with a suitable GUI aimed both at general purpose querying and extracting training data for AI models. Referring to the Fig.1, a MongoDB version of the data set was built using a semi-automated pipeline to create two kinds of JSON documents: the one collecting the information falling into the screenshots labeled as DESCRIPTION, and the one that gathers the information falling into the screenshots labeled both as CASE and TOPIC.
3.1 Data set extraction
Figure 2: Among the manually removed images there are presentation’s slide (a), manually (b) and automatic annotated images from DICOM viewer (c)
We decided to focus on a part of MedPix® that involves cases related with two diagnostic modalities, namely Computed Tomography (CT), and Magnetic Resonance Imaging (MRI). First of all, the images in the considered split were downloaded via Open-i®  [^5] https://openi.nlm.nih.gov/s: among them, there where some noisy images such as teaching materials or annotated images (Fig. 2) that are inconsistent as input image for training a LMM for which we expect a clear CT or MRI. A subsequent automatic scraping pipeline was implemented to extract the textual data related to the selected images using Selenium [^6] https://www.selenium.dev/ and Beautiful Soup [^7] https://www.crummy.com/software/BeautifulSoup/bs4/doc/. Finally, two kinds of JSON documents were devised to store, respectively, the information strictly connected with the images (descriptions document) and the one related to a clinical case (case-topic document). A one-to-many relation has been created between a clinical cases and images by embedding the U_id defined for a case-topic document in each descriptions document attached to each image related to the clinical case itself. An example of the two kinds of JSON documents is shown in Fig. 3.
Figure 3: An example of the JSON documents created in MedPix 2.0: (a) descriptions document, (b) case-topic document.
3.2 MongoDB representation
In order to process properly the data in MedPix 2.0 we built a MongoDB database to host all the JSON documents along with the images. The architectural choice stems not only from the nature of the data in MedPix 2.0 but also from considerations related to its high flexibility and scalability to distributed environments where also private multimodal medical data could be added to the original collections with the constraint of not being moved away from their generation site as it is the case of hospital generated information.
We built a MongoDB instance made by two collections, namely Image_ Descriptions, containing the descriptions documents, and Clinical_reports which contains all the case-topic documents. In our implementation, the images are stored in a separated folder, and they are accessed using a proper file:// URL built starting from their U_id. Finally, a view called Image_ Reports allow a direct access to both collections via their U_id.
A user-friendly GUI was built using PyQt5888https://www.riverbankcomputing.com/software/pyqt/ to allow an easy access to the database, querying it to obtain the desired data for visualization and/or download. As shown in Fig. 4.a, it is possible to select either the collection or the view to be queried and, add the query input in the relative fields. In Fig. 4.b, an example of the answer to a query is reported: samples matching the query are reported as a list of JSON objects and the user can save it or view a specific clinical case and/or image selected in list of the query results. An example of this view is shown in Fig. 5, where a MedPix® like GUI is reproduced to enforce usability for the users of the original website. In our GUI, the curated images scraped along with the texts are showed by default, but the user can choose to download the original (non curated) data related to the clinical case under investigation.
Figure 4: An example of the interface for querying the database where the user asks for CT scans of female patients (a) and the provided output (b).
Figure 5: An example of the view of a selected clinical case retrieved by the query posed in Fig. 4. The image, and the relative description are reported along with the information of the clinical case the image belongs to.
MedPix 2.0 and its querying interface can be a valid tool for both physicians and AI researchers since the desired data can be easily downloaded for further application, like training Deep Learning (DL) models where a large amount of structured data are required. The structured output downloaded from the interface, can be used in turn to fill a fixed template, and to provide a LMM with a rich textual prompt. Also multimodal tasks can be addressed if the textual descriptions are combined with the corresponding image, as we will demonstrate in Section 4. Our re-arranged version of MedPix®, contains the same information of the original dataset, but allows a easier usage for DL application in text-only, image-only and multimodal setup. This flexible structure is also suitable for adding new samples that meet the characteristics of the dataset, i.e. they should be a clear CT or MR scan with the associated textual data as in Fig. 3
4 Training a LMM using MedPix 2.0
The training-oriented reorganisation of MedPix 2.0 makes the data set suitable as training dataset for a multimodal deep neural network without further pre-processing of the data. To validate this claim, we used the obtained data set to build two classification tasks, namely scanning modality identification, i.e. determine if the given image is a CT scan or a MRI, and location classification, consisting in individuating the area of the body part shown. A balanced split from the data set was build via querying the database, and CLIP [16], that is one of the most recent and widespread multimodal models was used for both tasks. In particular, CLIP succeeds in achieving competitive performance in zero-shot contexts on a wide range of classification data sets by learning the relationships between images and their textual descriptions [20]. CLIP’s architecture consists of an image encoder and a text encoder and its training scheme involves the contrastive pre-training phase with a large corpus of image-text pairs, then the creation of the classifier from the text labels, and finally the zero-shot prediction.A schematic of this training is shown in Fig. 6.
Figure 6: The three characteristic phases of CLIP: (1) Contrastive pre-training, (2) Creation of the classifier from text labels and (3) Zero-shot prediction.
CLIP’s architecture is modular and adaptable to specific visual and textual encoders other than the default ones. Leveraging this peculiarity, we started a trial training phase to select the best suited visual encoder for our purposes, while keeping the textual encoder unchanged. Specifically, a ViT-L/14 and a RN50x16 were tested, and RN50x16, combined with CLIP’s default textual encoder, obtained an accuracy of 58% in the scanning modality identification task.
Once the visual encoder had been identified, we started testing a new textual encoder, BioBERT [9], in order to improve the overall classification performance. BioBERT has first been trained on the textual data obtained from MedPix 2.0 via the GUI, using the Mask Language Modelling train strategy, masking 15% of the input tokens, and then it was integrated with the RN50x16 encoder for a further CLIP fine-tuning on the selected train data. We trained both the models for 2500 epochs, with 30 as patience value, using a Cross Entropy loss and Adam [7] as optimizer with a learning rate of $1\mathrm{e}{-5}$. By doing so, the performance obtained increased considerably, reaching an accuracy of 88% in the scanning modality prediction task, and a 52,50% accuracy for the location classification task (Table 1). Prompt and labels details for both tasks are reported in Table 1.
Table 1: Overview of the two inference tasks reporting the used prompt, considered labels, and accuracy score
Task Scanning Modality classification Location classification
Prompt This image captured by {label} This image captured by CT or MR shows Disease located in the {label} area
Labels CT scan, Magnetic Resonance MR Head, Thorax, Adbomen, Spine and Muscles, Reproductive and Urinary System
Accuracy 88,00% 52,20%
5 Conclusions and Future work
In this paper we presented MedPix 2.0, a multimodal data set of clinical reports, CT and MR scans. We devised a semi-automated pipeline to download and curate the images in the original data sets, while structuring the textual information as a set of JSON documents collections which had been used to build a proper MongoDB instance. The NoSQL version of the data set can be accessed and queried with a usable GUI that has been developed purposely. Using the GUI one can browse the data set in the same manner as in the original MedPix® and can download the structured output of the query that is suitable for training AI models. To demonstrate this point, we developed a CLIP-based model for multimodal classification of CT/MR scans with respect to both the scanning modality and the body part shown in the input slice.
The presented data set and its MongoDB interface, represent in our view a relevant starting point for the development of AI multimodal models in the medical domain such as Information Extraction systems tailored for clinical reports, automated analysis of the medical images, or Generative AI models for clinical report generation as part of a Medical Decision Support System. All these systems can rely on MedPix 2.0 as a structured source of data containing both clinical cases and medical explanations about the found disease.
The structured JSON documents encode also implicit knowledge on the domain. We are currently working on a generative model that relies on a Knowledge Graph (KG) built from MedPix 2.0, to generate diagnostic findings given some preliminary information. The information retrieved analyzing the keys of the JSON documents can be easily structured as nodes (patients, diagnosis, treatment, exams, …) connected via edges symbolizing the relations be between them, and enriched by adding attributes at node or edge level. This re-organization of the data, can show some patterns on relations that otherwise cannot be easy recognized, and the developed graph, apart from AI-based applications, can be browsed by doctors for diagnostic or research purposes using a proper GUI. The scalability of the developed MongoDB database, makes it suitable for future extensions with the possibility to add brand new clinical cases or existing cases from data set mentioned in Section 2, that have to be compliant with privacy regulations and follow the required information structure. Moreover, the inherent distributed nature of MongoDB allows for creating huge databases across different wards where the data owned by a single institution do not need to be explicitly moved out of the hospital thus violating privacy regulations. The structure of MedPix 2.0 could also serve as a guide to develop suitable connectors to share allowed data in the EHDS. New cases can also be easily added to a KG that can be queried and used in a more complex pipeline with via a Retrieval Augmented Generation (RAG) approach to retrieve relevant piece of information for generating a more precise report via a generative LLM. Further improvement is being done also on the interface, providing the user with advanced data visualization tools, like the possibility to interactively compare similar cases, thus helping physicians during the diagnostic phase.
{credits}
5.1 Data and Code Availability
Source code and is freely available at https://github.com/CHILab1/MedPix-2.0.git. The data used for the test and train can be downloaded free of charge from the zenodo repository at the following link https://zenodo.org/records/12624810
5.1.1 Acknowledgements
This work is supported by the cup project B73C22000810001, project code ECS_00000022 “SAMOTHRACE” (Sicilian MicronanoTech Research And Innovation Center).
[…]
The authors declares that they have no relevant or material financial interests that relate to the research described in this paper.